# Knowledge Mechanisms in Large Language Models: A Survey and Perspective

### v2

@@ 6.4 How to Explore More Knowledge from Interdisciplinary Inspiration?
Neuroscience studies the structure and function of the brain at molecular, cellular, neural circuit, and neural network levels Squire et al. (2012). Generally, both mechanism analysis in LLMs and neuroscience utilize observation and intervention methods to investigate the basic principles of knowledge learning and memory, decision-making, language, perception, and consciousness. The biological signals of the human brain and the internal activation signals in LLMs are capable of reciprocal transformation Caucheteux et al. (2023); Feng et al. (2023a); Mossing et al. (2024); Flesher et al. (2021). Benefiting from advancements in neuroscience Jamali et al. (2024); de Schotten et al. (2022); Lee et al. (2022a), mechanism analysis in LLMs has identified analogous function neurons and regions Zhao et al. (2023a), and knowledge circuits Yao et al. (2024). Besides, leveraging plasticity theory in neuroscience, LLMs explain the underlying technical support for intelligence Sæbø and Brovold (2024). In the future, mechanism analysis of LLMs may draw inspirations from neuroscience, guiding the next generation of artificial intelligence in organizing neural frameworks and in the storage and utilization of knowledge Momeni et al. (2024); Yang et al. (2024b).
@@ 7.2 Embodied Intelligence
Recent studies Huang et al. (2024a); Chen et al. (2024b) have highlighted the discrepancy between the model’s performance in language tasks and its ability to integrate knowledge from different modalities effectively. These findings suggest that current models often prioritize linguistic information, failing to fully exploit the synergistic potential of multimodal data Wang et al. (2024c). There are some pioneering efforts in this direction  Pan et al. (2024a) , aiming to uncover the mechanisms by which multi-modal models store and retrieve information. Despite these advancements, there is still a need for further exploration to deepen our understanding of multi-modal knowledge storage.
@@
As discussed in the previous part, current language models are mainly based on tuning to gain knowledge, which requires a lot of training and high-quality data. These learnings are passive whereas, to be a human, evolution usually also undergoes communication and interaction. As an intelligent agent, the models should be able to learn through interactions and learn by themselves spontaneously. Recently, some work has attempted to enable the model to learn by themselves Zhang et al. (2024c) or learn by interaction with the environment Xu et al. (2024a); Xi et al. (2024). By integrating self-evolving mechanisms, models can continuously update their knowledge base and improve their understanding without relying solely on manually curated datasets. This not only reduces the dependency on large-scale labeled data but also allows the models to adapt to evolving linguistic norms and cultural contexts over time.
@@ 7.3 Domain LLMs
The success of general-purpose LLMs has indeed inspired the development of domain-specific models that are tailored to particular areas of knowledge, such as biomedicine Yu et al. (2024a); Moutakanni et al. (2024), finance Yang et al. (2023), geoscience Deng et al. (2023), ocean science Bi et al. (2024), etc. However, unlike human language, the knowledge of these different domains bears specific characteristics. It remains unclear whether LLMs can acquire complex scientific knowledge or if such knowledge still resides within the realm of current dark knowledge. Furthermore, does domain-specific knowledge such as mathematics share the same underlying mechanisms as textual knowledge Bengio and Malkin (2024), or does it exhibit more intricate mechanisms of knowledge acquisition? Currently, there is a relative lack of research focusing on the mechanism of these domain-specific knowledge and there is an increasing recognition of the importance of developing a deeper understanding of these mechanisms.
@@ 8 Conclusion
There are various forms of knowledge representation. However, due to current research constraints, this paper does not delve into space Li et al. (2024d), time Gurnee and Tegmark (2023), event-based knowledge, and geoscience Lin et al. (2024).
@@ B.2 Knowledge Mechanisms in Other Architectures
Surprisingly, similar mechanisms as those found in transformer-based LLMs have also been discovered in other architectural models. Specifically, Mamba employs the knowledge memorization mechanism similar to Transformer Sharma et al. (2024). Vision and multi-modal architectures also adopt function region (Modular Region Hypothesis) for knowledge utilization Pan et al. (2023a); Schwettmann et al. (2023); Koh et al. (2020); Bau et al. (2017), e.g., multi-modal neuron regions are responsible for multi-modal tasks. Besides, the connections hypothesis between neurons is found in vision architecture models Olah et al. (2020). Olah et al. (2020) further suggest that different types of knowledge reuse partial components, e.g., cars and cats reuse the same neurons (Reuse Hypothesis). As for the Dynamic Intelligence Hypothesis, it inherently focuses on entire artificial neural models.
@@ Appendix C Tools for Mechanism Analysis
Numerous tools exist for interpreting knowledge mechanisms in LLMs. TransformerLens Nanda and Bloom (2022) is a library for the mechanistic interpretability using observation and intervention. TransformerLens allows users to cache, remove, or replace internal activations during model running. XMD Lee et al. (2023a) provides various forms of feedback via an intuitiveness, which enable explanations align with the user feedback. NeuroX Dalvi et al. (2023) implements various interpretation methods under a unified API then provides interpretability of LLMs. PatchScope Ghandeharioun et al. (2024) is a tool developed by Google that employs a novel model to elucidate the hidden states in the original model. Transformer Debugger Mossing et al. (2024), an interpretability tool from OpenAI, utilizes GPT-4 and sparse auto-encoders to explain language neurons. Gao et al. (2024a) leverage sparse auto-encoders to extract interpretable features from a language model by reconstructing activations from a sparse bottleneck layer. Transcoders Dunefsky et al. (2024) decomposes model computations involving MLPs into interpretable circuits.

### v3

@@ 6.4 How to Explore More Knowledge from Interdisciplinary Inspiration?
Neuroscience studies the structure and function of the brain at molecular, cellular, neural circuit, and neural network levels Squire et al. (2012). Generally, both mechanism analysis in LLMs and neuroscience utilize observation and intervention methods to investigate the basic principles of knowledge learning and memory, decision-making, language, perception, and consciousness. The biological signals of the human brain and the internal activation signals in LLMs are capable of reciprocal transformation Caucheteux et al. (2023); Feng et al. (2023a); Mossing et al. (2024); Flesher et al. (2021). Benefiting from advancements in neuroscience Jamali et al. (2024); de Schotten et al. (2022); Lee et al. (2022a), mechanism analysis in LLMs has identified analogous function neurons and regions Zhao et al. (2023a), and knowledge circuits Yao et al. (2024). Besides, leveraging plasticity theory in neuroscience, LLMs explain the underlying technical support for intelligence Sæbø and Brovold (2024). In the future, mechanism analysis of LLMs may draw inspirations from neuroscience, guiding the next generation of artificial intelligence in organizing neural frameworks and in the storage and utilization of knowledge Ren and Xia (2024); Momeni et al. (2024); Yang et al. (2024b).
@@ 7.2 Embodied Intelligence
Recent studies Huang et al. (2024a); Chen et al. (2024b) have highlighted the discrepancy between the model’s performance in language tasks and its ability to integrate knowledge from different modalities effectively. These findings suggest that current models often prioritize linguistic information, failing to fully exploit the synergistic potential of multimodal data Wang et al. (2024c). There are some pioneering efforts in this direction  Pan et al. (2024a); Schwettmann et al. (2023a) , aiming to uncover the mechanisms by which multi-modal models store and retrieve information. Despite these advancements, there is still a need for further exploration to deepen our understanding of multi-modal knowledge storage.
@@
As discussed in the previous part, current language models are mainly based on tuning to gain knowledge, which requires a lot of training and high-quality data. These learnings are passive whereas, to be a human, evolution usually also undergoes communication and interaction. As an intelligent agent, the models should be able to learn through interactions and learn by themselves spontaneously. Recently, some work has attempted to enable the model to learn by themselves Zhang et al. (2024b) or learn by interaction with the environment Xu et al. (2024a); Xi et al. (2024). By integrating self-evolving mechanisms, models can continuously update their knowledge base and improve their understanding without relying solely on manually curated datasets. This not only reduces the dependency on large-scale labeled data but also allows the models to adapt to evolving linguistic norms and cultural contexts over time.
@@ 7.3 Domain LLMs
The success of general-purpose LLMs has indeed inspired the development of domain-specific models that are tailored to particular areas of knowledge Calderon and Reichart (2024), such as biomedicine Yu et al. (2024a); Moutakanni et al. (2024), finance Yang et al. (2023), geoscience Deng et al. (2023), ocean science Bi et al. (2024), etc. However, unlike human language, the knowledge of these different domains bears specific characteristics. It remains unclear whether LLMs can acquire complex scientific knowledge or if such knowledge still resides within the realm of current dark knowledge. Furthermore, does domain-specific knowledge such as mathematics share the same underlying mechanisms as textual knowledge Bengio and Malkin (2024), or does it exhibit more intricate mechanisms of knowledge acquisition? Currently, there is a relative lack of research focusing on the mechanism of these domain-specific knowledge and there is an increasing recognition of the importance of developing a deeper understanding of these mechanisms.
@@ 8 Conclusion
There are various forms of knowledge representation. However, due to current research constraints, this paper does not delve into space Li et al. (2024f), time Gurnee and Tegmark (2023), event-based knowledge, and geoscience Lin et al. (2024).
@@
Acknowledgements
We would like to express gratitude to the anonymous reviewers for their kind comments. This work was supported by the National Natural Science Foundation of China (No. 62206246, No. NSFCU23B2055, No. NSFCU19B2027), the Fundamental Research Funds for the Central Universities (226-2023-00138), Zhejiang Provincial Natural Science Foundation of China (No. LGG22F030011), Yongjiang Talent Introduction Programme (2021A-156-G), Information Technology Center and State Key Lab of CAD&CG, Zhejiang University, and NUS-NCS Joint Laboratory (A-0008542-00-00).
@@ B.2 Knowledge Mechanisms in Other Architectures
Surprisingly, similar mechanisms as those found in transformer-based LLMs have also been discovered in other architectural models. Specifically, Mamba employs the knowledge memorization mechanism similar to Transformer Sharma et al. (2024). Vision and multi-modal architectures also adopt function region (Modular Region Hypothesis) for knowledge utilization Pan et al. (2023a); Schwettmann et al. (2023a); Koh et al. (2020); Bau et al. (2017), e.g., multi-modal neuron regions are responsible for multi-modal tasks. Besides, the connections hypothesis between neurons is found in vision architecture models Olah et al. (2020). Olah et al. (2020) further suggest that different types of knowledge reuse partial components, e.g., cars and cats reuse the same neurons (Reuse Hypothesis). As for the Dynamic Intelligence Hypothesis, it inherently focuses on entire artificial neural models. Generally, neural models across various architectures, trained with different objectives on different data and modalities, are converging to a shared statistical model of reality in their representation spaces Huh et al. (2024). These neural models may tend to share similar knowledge mechanisms and imagination Zhou et al. (2024)
@@ Appendix C Tools for Mechanism Analysis
Numerous tools exist for interpreting knowledge mechanisms in LLMs. TransformerLens Nanda and Bloom (2022) is a library for the mechanistic interpretability using observation and intervention. TransformerLens allows users to cache, remove, or replace internal activations during model running. XMD Lee et al. (2023a) provides various forms of feedback via an intuitiveness, which enable explanations align with the user feedback. NeuroX Dalvi et al. (2023) implements various interpretation methods under a unified API then provides interpretability of LLMs. PatchScope Ghandeharioun et al. (2024) is a tool developed by Google that employs a novel model to elucidate the hidden states in the original model. Transformer Debugger Mossing et al. (2024), an interpretability tool from OpenAI, utilizes GPT-4 and sparse auto-encoders to explain language neurons. Sparse autoencoders Gao et al. (2024b) leverages sparse auto-encoders to extract interpretable features from a language model by reconstructing activations from a sparse bottleneck layer. Transcoders Dunefsky et al. (2024) decomposes model computations involving MLPs into interpretable circuits.